# LLaVA-3D: A Simple yet Effective Pathway to Empowering LMMs with 3D-awareness

### v1

@@ 4 Training
To leverage the 2D priors from established 2D LMMs, we train our LLaVA-3D model based on the pre-trained LLaVA-1.5. Considering the scarcity of 3D scene-language data, our training procedure comprises two stages, each focusing on different training targets of the model.
Stage 1: 3D Patch Language Alignment.
During the first training stage, we use the region-level and scene-level caption data that describe spatial relationships among 3D objects to align the 3D patches with the LLM for enhanced 3D spatial comprehension. At this stage, the input multi-view images used are selected from sequences that correspond to specific regions or entire scenes. We freeze the vision encoder and LLM parameters, and only train the projection layer and 3D position embedding layer, encouraging efficient alignment between 3D patch features and text space. Since 3D patches are derived from CLIP features augmented with 3D positional information, the alignment between 3D Patch and LLM converges rapidly.
Figure 3: LLaVA-3D-Instruct-1M. The hybrid 2D and 3D Dataset Collection. Left: Distribution of data across categories, with the outer circle representing all categories and the inner circle illustrating data subset distribution. Right: Detailed dataset quantities.
Stage 2: Task Instruction Tuning.
During the instruction-tuning stage, LLaVA-3D is optimized to respond to complex 3D V&L tasks and maintain its inherent 2D image reasoning and instruction-following capabilities. To facilitate this capability, we collect the LLaVA-3D-Instruct-1M dataset, a hybrid collection of 2D and 3D data specifically tailored for instruction tuning. The overall distribution of the dataset collection is shown in Fig 3, more details about the instructional tuning datasets are listed in the appendix. The 2D Data of LLaVA-3D-Instruct-1M is derived from existing LLaVA-1.5 instruction tuning data, ensuring the preservation of 2D image comprehension and vision-language conversation abilities. When tuning with 2D data, we keep the vision encoder frozen and jointly train the projection layer and LLM. The 3D Data of LLaVA-3D-Instruct-1M, on the other hand, comprises data from diverse 3D QA, 3D dense captioning, and 3D grounding tasks. During the 3D data instruction tuning, the 3D position embedding layer will be added to jointly train with the other modules. Additionally, for tasks where the instruction contains 3D coordinate information or requires 3D bounding box outputs, the corresponding encoding and decoding modules will be trained together. During instruction tuning, the 3D data pathway includes the 3D position embeddings and 3D patches, while the 2D data pathway is the original LLaVA. All modules except for the 3D position embeddings to construct 3D patches are shared across 2D and 3D data. This training setup ensures that LLaVA-3D is capable of processing both 2D and 3D visual tokens effectively, and is adaptive to various task formulations and requirements.
@@ 5 Experiments
Table 1: Quantitative comparison with SOTA models on various 3D QA tasks. “C” stands for “CIDEr”, “B-4” for “BLEU-4”, “M” for “METEOR”, “R” for “ROUGE”, “Sim” for sentence similarity, and “EM@1” for top-1 exact match. Gray indicates evaluation results with refined exact-match protocol.
@@ C B-4 M R EM@1 EM@1
Task-specific fine-tuned 3D LMMs
@@
LLaVA-3D 91.7 14.5 20.7 50.1 27.0 (45.0) 55.6 (57.6)
Table 2: Quantitative comparison on MMScan QA benchmark. “S.-BERT", “B-1”, “B-4”, “R.-L.”, “MET.” represents “Sentence-BERT", “BLEU-1”, “BLEU-4”, “ROUGE-L”, “METEOR”, respectively. Here, we report the top-1 exact match with (the refined exact-match protocol results) for “EM@1”.
@@
LLaVA-3D Generalist 52.3 61.2 54.4 28.7 61.2 44.5 43.6 74.6 76.3 38.7 13.1 55.5 19.5 45.2 (51.4)
Table 3: Quantitative comparison with SOTA models on OpenEQA benchmark
@@
LLaVA-3D 20 51.2
In this section, we conduct extensive evaluations to examine the capabilities of LLaVA-3D. To begin with, we introduce the implementation details (Sec. 5.1). Then, we compare our model’s 3D scene understanding (Sec. 5.25.35.4)and 2D image understanding (Sec. 5.5) capability with previous methods. Finally, we conduct a thorough analysis to validate the effectiveness of the components and designs of our LLaVA-3D (Sec. 5.65.7).
5.1 Implementation Details
LLaVA-3D is built upon the LLaVA-1.5-7B, utilizing their pre-trained weights from the HuggingFace library. For 3D tasks, we add the 3D position embeddings to the 2D patch visual tokens, and utilize the voxelization pooling strategy to reduce 3D patch number before passing the input visual tokens to the projection layer and LLM. The number of views $V$ is set to be 20 and voxel size is set to 0.2m. Due to the LLM context length limitation, the maximum number of 3D patch tokens after 3D-aware pooling is set to 3096. For 2D tasks, LLaVA-3D functions the same as LLaVA. All experiments are conducted on 8 × 80G A100 GPUs. We train our model for 1 epoch with a learning rate of 1e-3 and a batch size of 32 in stage 1, and fine-tune on the collected LLaVA-3D-Instruct-1M dataset, with a learning rate of 2e-5 and a batch size of 16 in stage 2.
@@ 5.2 Evaluation on 3D Question Answering
3D Question Answering requires a model to generate responses to the natural language queries questioning towards a 3D scene. In this section, we validate LLaVA-3D performance on various 3D question answering benchmarks: ScanQA [3], SQA3D [40], MMScan QA [39], and OpenEQA [41].
Spatial Understanding with ScanQA and SQA3D.
@@
ScanQA and SQA3D are both built on ScanNet dataset. The ScanQA dataset consists of 41363 questions about 800 scenes, including 32337 unique questions. Its validation set contains 4675 questions about 71 scenes. SQA3D comprises 20.4k descriptions of 6.8k unique situations collected from 650 ScanNet scenes and 33.4k questions about these situations. Questions in ScanQA require basic recognition and 3D reasoning capabilities, and SQA3D further incorporates the situation understanding and situated reasoning into embodied 3D scene understanding. Following prior works, we adopt BLEU scores, METEOR, ROUHE-L, CIDEr and EM (“exact match”) as our evaluation metrics for ScanQA and EM for SQA3D respectively. As shown in Tab. 1, current 2D LMMs fail to achieve competitive performance with the latest 3D LMMs trained with 3D awareness, which might be attributed to the lack of explicit 3D representation. Besides, compared with task-specific fine-tuned 3D LMMs that need to be further fine-tuned on the corresponding datasets, our LLaVA-3D could perform as a generalist and achieve the SOTA performance on these benchmarks.
Coordinate Spatial Understanding with MMScan QA.
@@
MMScan QA includes 5.2k scans from ScanNet, 3RScan, and Matterport3D, along with 116k training questions and 29k validation questions. These questions span existential inquiries, attribute understanding, and more advanced queries. Unlike ScanQA and SQA3D, some MMScan QA questions require 3D reasoning based on object coordinates, rather than relying solely on text descriptions, demanding the model capable of understanding 3D coordinates information. All data samples are classified into one of the following sub-categories: ST-attr, ST-space, OO-attr, OO-space, OR, where ST stands for Single-target, attr for attribute, OO for Object-Object, and OR for Object Region. Besides, there is a minor part of QA samples for advanced understanding and reasoning, such as situated QA related to everyday life. We present the results under GPT-4 evaluation, data-driven metrics, and traditional metrics respectively in Tab. 2. In this benchmark, the well-trained LL3DA and LEO are further fine-tuned on the full 1.2M MMScan QA training dataset. Our LLaVA-3D, trained on LLaVA-3D-Instruct-1M (which includes 440K MMScan QA training samples), achieves significantly better performance on the MMScan QA benchmark compared to the specially fine-tuned LL3DA and LEO. The results highlight the training efficiency of LLaVA-3D and its strong 3D understanding ability to serve as the generalist model.
Embodied Question Answering with OpenEQA.
@@
OpenEQA is the first open-vocabulary benchmark designed for spatial understanding and embodied reasoning in embodied question answering, specifically in the era of foundation models. It features an automated evaluation protocol powered by LLMs, which shows strong alignment with human judgment. Our evaluations are conducted using the EM-EQA data split of OpenEQA, which includes over 1,600 high-quality, human-generated questions from diverse real-world environments. In EM-EQA, the model is required to generate a textual answer to a question based on an episode history, which typically consists of RGB images, depth information, camera poses, and intrinsic camera data. The results in Tab. 3 demonstrate that LLaVA-3D surpasses Claude3 and Gemini-Pro, and achieves comparable performance with powerful GPT-4V on this benchmark with significantly fewer model parameters.
@@
Table 4: Quantitative Comparisons with SOTA models for 3D Dense Captioning on Scan2Cap. “C” stands for “CIDEr”, “B-4” for “BLEU4”, “M” for “METEOR”, “R” for “ROUGE”, “Sim” for sentence similarity, and “EM@1” for top-1 exact match. The n-gram metrics for Scan2Cap are governed by IoU@0.5.
@@
Scan2Cap [13] 39.08 23.32 21.97 44.78
MORE [24] 40.94 22.93 21.66 44.42
SpaCap3D [46] 44.02 25.26 22.33 45.36
D3Net [7] 46.07 30.29 24.35 51.67
UniT3D [14] 46.69 27.22 21.91 45.98
3DJCG [5] 49.48 31.03 24.22 50.80
3D-VLP [25] 54.94 32.31 24.83 51.51
3D-VisTA [52] 61.60 34.10 26.80 55.00
Vote2Cap-DETR [11] 61.81 34.46 26.22 54.40
LL3DA [12] 65.19 36.79 25.97 55.06
LEO [20] 72.4 38.2 27.9 58.1
LLaVA-3D 79.21 41.12 30.21 63.41
@@ 5.3 Evaluation on 3D Dense Captioning
3D dense captioning requires the model to localize all the objects in a 3D scene and then generate descriptive sentences for each object. To evaluate our model on the 3D dense captioning tasks, we utilize the off-the-shelf segmentation model Mask3D to generate object proposals. Then we further construct the 3D coordinate tokens based on the 3D object center coordinates to guide the model to perform the task. Additionally, we utilize two types of textual instructions that prompt the model to either describe the object or describe and output the bounding box of the object. We report the performance of various methods on two 3D dense captioning benchmarks: Scan2Cap and MMScan Captioning.
Scan2Cap.
Scan2Cap requires the model to describe the object’s appearance and the spatial relations with nearby objects and output the corresponding 3D bounding box. As illustrated in Tab. 4, our method consistently outperforms the existing method on the Scan2Cap benchmark.
MMScan Captioning.
MMScan Captioning focuses on identifying common aspects of 3D objects such as Object Type, Color, Shape, Position, Function, and Design. The benchmark utilizes GPT4 to assess whether these aspects are correct in the object description. We benchmark various methods on the MMScan Captioning benchmark in Tab. 5. The results show that our method surpasses existing approaches across all metrics by a substantial margin, especially achieving a 49.5% improvement in the Color score and a 43.3% improvement in the Design score. This significant performance boost can be attributed to our 2D LMM-based architecture.
Uniquely, LLaVA-3D takes multi-view images as inputs, enabling a user-friendly feature where users can simply click on selected images to generate both 3D object captions and 3D bounding boxes. We refer to this capability as 2D Click-based 3D Dense Captioning. (as illustrated in Fig. 4).
@@ 5.4 Evaluation on 3D Visual Grounding
3D visual grounding aims to localize the target object in the 3D scene using natural language descriptions. In this section, we initially report the performance on the ScanRefer [6] benchmark. For quantitative comparisons, we include both task-specific approaches and generalist models: the state-of-the-art specialists in 3D VG and generalist models like LLM-Grounder [48], 3D-LLM [18], and Chat3D-v2 [19]. 3D-LLM [18]] uses the location tokens to discretize the 3D scene and predicts the bounding box as the location tokens that were added to the vocabulary. Chat-3D v2 [19] first detects all the objects and then identifies the object that best matches the provided description. The results in Tab. 6 demonstrate that by combining with the grounding module, our LLaVA-3D could achieve even better performance compared with the task-specific specialists model as a generalist. Such a paradigm allows LLaVA-3D to integrate with various powerful 3D grounding modules and inject the 3D understanding capability and world knowledge into the grounding process, as illustrated in Fig. 7.
Table 6: Quantitative comparison with SOTA models on various 3D VG tasks.
ScanRefer
Acc@0.25 Acc@0.5
@@
ScanRefer [6] 37.3 24.3
MVT [21] 40.8 33.3
3DVG-Trans [50] 45.9 34.5
ViL3DRel [10] 47.9 37.7
BUTD-DETR [22] 52.2 39.8
ReGround3D [51] 53.1 41.1
Task-specific fine-tuned 3D LLMs
LLM-Grounder [48] 17.1 5.3
3D-LLM [18] 30.3 -
Chat3D-v2 [19] 35.9 30.4
LLaVA-3D 54.1 42.2
@@
Table 7: Quantitative Comparisons with SOTA models on zero-shot 2D benchmarks.
[…]
MobileVLM [15] MLLaMA 2.7B 336 47.5 59.6 1289 -
InstructBLIP [16] Vicuna-7B 224 50.1 36.0 26.2
InstructBLIP [16] Vicuna-13B 224 50.7 1213 25.6
Qwen-VL [4] Qwen-7B 448 63.8 38.2
Qwen-VL-Chat [4] Qwen-7B 448 61.5 60.6 1488
Shikra [9] Vicuna-13B 224 58.8
LLaMA-VID [33] Vicuna-7B 336 65.1 1521
LLaMA-VID [33] Vicuna-13B 336 66.6 1542
LLaVA-1.5 [38] Vicuna-7B 336 58.2 65.2 1511 31.1
LLaVA-1.5 [38] Vicuna-13B 336 61.3 69.2 1531/295 36.1
LLaVA-3D Vicuna-7B 336 57.8 65.0 1502 30.9
@@ 5.5 Evaluation on 2D benchmarks
Since our model is trained on a mix of 2D and 3D datasets, we evaluate it across various 2D benchmarks to ensure it retains the 2D image understanding capabilities of the original LLaVA. As demonstrated in Tab. 7, LLaVA-3D achieves performance comparable to that of LLaVA-1.5 across several 2D image understanding benchmarks. This performance underscores the architectural superiority of our model compared to other 3D LMMs.
Table 8: Architecture Comparison on various 3D V&L Benchmark.
3D Feature Connector LLM / LMM ScanQA SQA3D Inference time
(a) (SAM + CLIP) w / PE Q-Former Vicuna-7B 21.9 49.3 900s
(b) (SAM + CLIP) w / PE Pooling + MLP Vicuna-7B 22.1 49.2 900s
(c) CLIP w / PE Pooling + MLP Vicuna-7B 23.4 51.2 0.2s
(d) CLIP w / PE Pooling + MLP LLaVA-1.5 27.0 55.6 0.2s
(e) CLIP w / PE Pooling + MLP PLLaVA 27.9 56.2 0.2s
Table 9: Effectiveness of 3D Patch Representation. Training using the instruction tuning datasets, the only difference between multi-image LLaVA and LLaVA-3D is the patch type.
Method Patch Type ScanQA SQA3D MMScan QA Scan2Cap
multi-image LLaVA 2D 24.1 52.3 32.7 29.1
LLaVA-3D 3D 27.0 (+2.9) 55.6 (+3.3) 41.5 (+8.8) 79.2 (+50.1)
5.6 Architecture Analysis
In this section, we delve deeper into the architectural benefits and efficacy of adapting a 2D large multimodal model (LMM) to a 3D LMM, as opposed to developing a 3D LMM solely from LLMs. A notable approach in the latter category is 3D-LLM, which leverages foundational 2D visual models such as SAM and CLIP for feature extraction. This method then employs 3D multi-view fusion to generate 3D point features, followed by the use of a Q-Former to condense these point features into a fixed number of tokens.
Architecture Comparison.
To ensure the fairness of the experiment as much as possible, starting from LLM, we first ablate different 3D feature encoders, and 3D-language connectors using the same instruction tuning datasets. As shown in Tab. 8, the Q-Former (a) and Pooling + MLP (b) share a similar performance on 3D V&L benchmarks. Notably, using CLIP (c) alone instead of SAM + CLIP (b) achieves better performance and significantly reduces 3D feature extraction computation time from 900s to 0.2s.
Effectiveness of 3D Patch.
Once we don’t decorate the 2D patches with corresponding 3D position embedding, LLaVA-3D will degenerate into 2D multi-image LMMs. To further ascertain the efficacy of our proposed 3D representation: 3D Patch, we conduct additional experiments across a variety of 3D question answering and 3D dense captioning benchmarks. As shown in Tab. 9, integrating 2D patches within a 3D context enhances the model’s 3D spatial understanding and reasoning capabilities, resulting in 2.9% and 3.3% improvements on the ScanQA and SQA3D benchmarks, respectively. Additionally, 3D patches are crucial for tasks that require explicit 3D world modeling, leading to significant improvements of 8.8% on the MMScan QA benchmark and an impressive 50.1% on the Scan2Cap benchmark.
Benefits from pre-trained 2D LMM.
Leveraging our foundation in 2D LMMs, our framework benefits significantly from the robust pre-training on extensive image/video-text datasets. In Tab. 8, we explore the advantages of initializing from a pre-trained 2D LMM compared to starting directly from an LLM. The experimental results consistently demonstrate that starting with a well-tuned 2D LMM substantially could enhance performance in 3D spatial understanding tasks. Besides, we observe that stronger base 2D LMM could lead to better 3D spatial understanding performance. Besides, we find that initializing from a pre-trained 2D LMM could achieve 3.5× faster training convergence speed of LLaVA-3D compared with existing 3D LMMs [12, 20]. More details could be found in our appendix.

### v3

@@ 4 Training
To achieve strong 3D understanding and grounding capabilities without compromising 2D understanding, we conduct the two-stage training strategy, inspired by  [49]. The first stage equips the model with various 2D and 3D task instruction following capabilites, and the second stage further enhances the 3D grounding capability.
Stage 1: Multi-Task Instruction Tuning. During the instruction-tuning stage, LLaVA-3D is optimized to respond to complex 3D V&L tasks while maintaining its original 2D reasoning and instruction-following capabilities. To facilitate this capability, we fine-tune the model on the joint 2D and 3D data. Specifically, the 2D data is sampled from the LLaVA-Video training dataset, ensuring the preservation of 2D video comprehension and conversation abilities. For the 3D data, we collect the LLaVA-3D-Instruct-86K dataset, a hybrid collection of 3D data specifically tailored for instruction tuning covering various 3D tasks. The overall distribution of the dataset collection is shown in Fig 3. For the 3D tasks, the 3D position encoding layer will be added to jointly train with the other modules. Additionally, for tasks which requires 3D bounding box outputs, the grounding decoder will be trained together. This training setup ensures that LLaVA-3D can effectively process both 2D and 3D visual tokens and is adaptive to various tasks.
Figure 3: LLaVA-3D-Instruct-86K. The 3D Dataset Collection. Left: Distribution of data across categories, with the inner circle representing all categories and the outer circle illustrating data subset distribution. Right: Detailed dataset quantities.
Stage 2: Decoder-only Fine-tuning. Since the grounding decoder fails to converge within a single epoch during the first stage of training, we further train the grounding decoder for additional epochs using the 3D visual grounding data while keeping all other components frozen. During this stage, the location token is further trained with the grounding decoder for better 3D visual grounding performance, without influencing the strong 2D and 3D scene understanding capabilities.
@@ 5 Experiments
Table 1: Quantitative comparison with SOTA models on various 3D QA tasks. “C” stands for “CIDEr”, “B-4” for “BLEU-4”, “M” for “METEOR”, “R” for “ROUGE”, and “EM@1” for top-1 exact match. Gray indicates evaluation results with refined exact-match protocol.
@@ C B-4 M R EM@1 EM@1
3D LMMs
@@
LLaVA-3D 103.1 16.4 20.8 49.6 30.6 60.1
Table 2: Quantitative comparison on MMScan QA benchmark. “ST” stands for Single-target, “attr” for attribute, “OO” for Object-Object, and “OR” for Object Region.“S.-BERT”, “B-1”, “B-4”, “R.-L.”, “MET.” represents “Sentence-BERT", “BLEU-1”, “BLEU-4”, “ROUGE-L”, “METEOR”, respectively. Here, we report the top-1 exact match with (the refined exact-match protocol results) for “EM@1”.
@@
LLaVA-3D Generalist 55.4 63.2 57.1 34.1 63.2 47.5 44.9 76.2 78.3 39.2 13.9 57.5 20.3 50.1 (54.9)
Table 3: Quantitative comparison with SOTA models on OpenEQA benchmark.
@@
LLaVA-3D 32 53.2
In this section, we conduct extensive evaluations to examine the capabilities of LLaVA-3D, based on LLaVA-Video. We compare our model’s 3D scene understanding (Sec. 5.15.25.3) and 2D video understanding (Sec. 5.4) capability with previous methods. Then we thoroughly analyze the effectiveness of the components and designs of LLaVA-3D and demonstrate the our LLaVA-3D can be integrated with any 2D LMMs to equip 3D capabilities (Sec. 5.5).
@@ 5.1 Evaluation on 3D Question Answering
3D Question Answering requires a model to generate responses to the natural language questions in a 3D scene. In this section, we validate LLaVA-3D performance on various 3D question answering benchmarks: ScanQA [3], SQA3D [39], MMScan QA [38], and OpenEQA [40].
@@
Spatial Understanding with ScanQA and SQA3D. ScanQA and SQA3D are both built on the ScanNet dataset. The ScanQA dataset consists of 41363 questions about 800 scenes, including 32337 unique questions. SQA3D comprises 20.4k descriptions of 6.8k unique situations collected from 650 ScanNet scenes and 33.4k questions about these situations. Questions in ScanQA require basic recognition and 3D reasoning capabilities, and SQA3D further incorporates situation understanding and situated reasoning into embodied 3D scene understanding. As shown in Tab. 1, LLaVA-Video could achieve promising performance on these benchmarks in a zero-shot manner, even surpassing the task-specific methods. This phenomenon may suggest that these benchmarks do not truly assess the model’s 3D spatial understanding ability. Notably, our model could achieve the SOTA performance on these benchmarks.
@@
Coordinate Spatial Understanding with MMScan QA. MMScan QA includes 5.2k scans from ScanNet, 3RScan, and Matterport3D, along with 116k training questions and 29k validation questions. These questions span existential inquiries, attribute understanding, and more advanced queries. Unlike ScanQA and SQA3D, some MMScan QA questions require 3D reasoning based on object coordinates rather than relying solely on text descriptions, demanding the model capable of understanding 3D coordinates information. We present the results under GPT-4 evaluation, data-driven metrics, and traditional metrics respectively in Tab. 2. Our LLaVA-3D achieves significantly better performance compared to LL3DA and LEO which are further fine-tuned on full MMScan QA. The results highlight the training efficiency of LLaVA-3D and its strong 3D understanding ability to serve as the generalist model.
@@
Embodied Question Answering with OpenEQA. OpenEQA is the first open-vocabulary benchmark designed for spatial understanding and embodied reasoning in embodied question answering. It features an automated evaluation protocol powered by LLMs, which shows strong alignment with human judgment. Our evaluations are conducted using the EM-EQA data split of OpenEQA, which includes over 1,600 high-quality, human-generated questions from diverse real-world environments. The results in Tab. 3 demonstrate that LLaVA-3D surpasses Claude3 and Gemini-Pro, and achieves comparable performance with powerful GPT-4V on this benchmark with significantly fewer model parameters.
@@
Table 4: Quantitative Comparisons with SOTA models for 3D Dense Captioning on Scan2Cap. The n-gram metrics for Scan2Cap are governed by IoU@0.5.
@@
Scan2Cap [10] 39.1 23.3 22.0 44.8
3D-VLP [23] 55.0 32.3 24.8 51.5
3D-VisTA [57] 61.6 34.1 26.8 55.0
Vote2Cap-DETR [7] 61.8 34.5 26.2 54.4
LL3DA [8] 65.2 36.8 26.0 55.0
LEO [18] 68.4 36.9 27.7 57.8
ChatScene [17] 77.2 36.3 28.0 58.1
LLaVA-3D 84.1 42.6 29.0 63.4
@@ 5.2 Evaluation on 3D Dense Captioning
3D dense captioning requires the model to localize all the objects in a 3D scene and then generate a descriptive sentence for each object. To evaluate our model on the 3D dense captioning tasks, we utilize the off-the-shelf 3D instance segmentation model [44] to generate object proposals. Then we further construct the 3D coordinate tokens based on the 3D object center coordinates to guide the model to handle the task. We report the performance of various methods on two 3D dense captioning benchmarks:
Scan2Cap. Scan2Cap requires the model to describe the object’s appearance and the spatial relations with nearby objects and output the corresponding 3D bounding box. As illustrated in Tab. 4, our method consistently outperforms the existing method on the Scan2Cap benchmark.
MMScan Captioning. MMScan Captioning focuses on identifying common aspects of 3D objects such as Object Type, Color, Shape, Position, Function, and Design. We benchmark various methods on MMScan Captioning benchmark in Tab. 5. The results show that our method surpasses existing approaches across all metrics by a substantial margin, especially achieving a 49.5% and 43.3% improvement in the Color score and the Design score respectively. The strong performance further demonstrates the advantages of architectures based on 2D LMMs.
Uniquely, LLaVA-3D takes multi-view images as inputs, enabling a user-friendly feature where users can simply click on the selected images to generate both 3D object captions and 3D bounding boxes, as illustrated in Fig. 4.
@@ 5.3 Evaluation on 3D Visual Grounding
3D visual grounding aims to localize the target object in the 3D scene using the natural language description. In this section, we initially report the performance on the ScanRefer [4] and Multi3DRefer [52] benchmarks in Tab. 6. Previous task-specific methods [57, 20, 56] typically rely on point clouds extracted from 3D reconstructed meshes as input, which are not easily accessible in real-world applications. Meanwhile, current 3D LMMs [17, 16, 9] tend to decouple the grounding task into two stages: first, extract the 3D objects from the 3D scene using the off-the-shelf 3D segmentor and then convert the task into a 3D object selection task. Without relying on the constructed point cloud, our method could directly decode the accurate 3D bounding boxes from 3D patches and achieve the SOTA performance (49.8 Acc@0.25 on Multi3DRefer) in the single-stage manner.
Table 6: Quantitative comparison with SOTA models on various 3D VG tasks. $\dagger$ represents that we apply the current two stage 3D visual grounding approach to LLaVA-3D.
ScanRefer Multi3DRefer
Acc@0.25 Acc@0.5 Acc@0.25 Acc@0.5
@@
ScanRefer [4] 37.3 24.3 - -
MVT [19] 40.8 33.3 - -
3DVG-Trans [54] 45.9 34.5 - -
ViL3DRel [6] 47.9 37.7 - -
BUTD-DETR [20] 52.2 39.8 - -
ReGround3D [56] 53.1 41.1 - -
M3DRef-CLIP [56] 51.0 44.7 42.8 38.4
Two-Stage 3D LMMs
Chat-3D v2 [16] 35.9 30.4 - -
Grounded 3D-LLM [9] 47.9 44.1 45.2 40.6
Chat-Scene [17] 55.5 50.2 57.1 52.4
LLaVA-3D 63.9 58.6 68.1 62.9
Single-Stage 3D LMMs
3D-LLM [15] 30.3 - - -
LLaVA-3D 50.1 42.7 49.8 43.6
@@
Table 7: Quantitative Comparisons on 2D video benchmarks.
Method MVBench VideoMME
LLaVA-Video [53] 58.6 63.3
LLaVA-3D 58.1 62.8
@@ 5.4 Evaluation on 2D benchmarks
Since our model is trained on the joint 2D and 3D datasets, we evaluate it on the 2D video benchmarks to ensure it retains the 2D understanding capabilities of the original LLaVA-Video. As demonstrated in Tab. 7, LLaVA-3D achieves performance comparable to LLaVA-Video across various 2D video understanding and conversation benchmarks, which the current existing 3D LMMs do not possess. This performance highlights the architectural advantages of our model over other 3D LMMs.
Table 8: Analysis of model architecture and generalization on various 3D V&L Benchmark.
3D Feature Connector LLM / LMM ScanQA SQA3D ScanRefer Inference time
(a) (SAM + CLIP) w / PE Q-Former Vicuna-7B 21.9 49.3 - 900s
(b) (SAM + CLIP) w / PE Pooling + MLP Vicuna-7B 22.1 49.2 - 900s
(c) CLIP w / PE Pooling + MLP Vicuna-7B 23.4 51.2 43.8 0.2s
(d) CLIP w / PE Pooling + MLP LLaVA-1.5-7B [34] 27.0 55.6 47.9 0.2s
(e) CLIP w / PE MLP InternVL2.5-7B [12] 29.1 58.8 49.3 0.3s
(f) CLIP w / PE MLP LLaVA-Video-7B [53] 30.6 60.1 50.1 0.2s
Table 9: Effectiveness of 3D Patch Representation
Patch Type ScanQA SQA3D MMScan QA Scan2Cap
2D 29.4 59.8 42.1 29.7
3D 29.8 (+0.4) 60.1 (+0.3) 55.4 (+13.3) 84.1 (+54.4)